Skip to main contentSkip to solution

Let A and B be two events such that P(A)=0.25, P(B)=0.50 and P(A ∩ B)=0.14. The probability that neither A nor B occurs is:

Solution

✅ Correct Option: 3

Given:

P(A)=0.25P(A) = 0.25

P(B)=0.50P(B) = 0.50

P(A∩B)=0.14P(A ∩ B) = 0.14

To find P(neither A nor B) =P(A′∩B′)=1−P(A∪B)= P(A' ∩ B') = 1 - P(A ∪ B)

Using: P(A∪B)=P(A)+P(B)−P(A∩B)P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

=0.25+0.50−0.14= 0.25 + 0.50 - 0.14

=0.61= 0.61

Therefore, P(neither A nor B) =1−0.61=0.39= 1 - 0.61 = 0.39

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question